1.) calculate the heat energy required to raise the temperature of : i) 1.0kg water from 0degree Celsius to 100 degrees Celsius. ii) 10kg of copper from 20 degrees to 30degrees Celsius. iii) 6.0kg of aluminum from 20 degrees to 60 degrees. iv) 4.0kg of mercury from 10 degrees to 70degrees.
1
Expert's answer
2021-10-24T18:25:26-0400

i)

"Q=mc\Delta T=1.0\ kg\times4200\ \dfrac{J}{kg\!\times^{\circ}\!C}\cdot100^{\circ}C=420000\ J."

ii)

"Q=mc\Delta T=10\ kg\times385\ \dfrac{J}{kg\!\times^{\circ}\!C}\cdot10^{\circ}C=38500\ J."

iii)

"Q=mc\Delta T=6.0\ kg\times900\ \dfrac{J}{kg\!\times^{\circ}\!C}\cdot40^{\circ}C=216000\ J."

iv)

"Q=mc\Delta T=4.0\ kg\times138\ \dfrac{J}{kg\!\times^{\circ}\!C}\cdot60^{\circ}C=33120\ J."